Find the slope and y-intercept of the line. Write fractions in lowest form.
The slope is what?
The y-intercept is what?
4y+5x+4=15+5x

Answer by stanbon(75887)

The slope is what?
The y-intercept is what?
4y+5x+4=15+5x
Subtract 5x from both sides to get:
4y+4 = 15
4y = 11
y = 11/4
-------------
slope = 0
y-intercept = 11/4
=======================
Cheers,
Stan H.